Use a fraction wall to compare the following fractions.

a. 1/3 [ ] 1/4

b. 4/7 [ ] 2/5

c. 2/8 [ ] 1/4

d. 2/5 [ ] 1/2

e. 4/8 [ ] 3/4

f. 4/5 [ ] 1/1


1
Expert's answer
2021-06-14T15:10:20-0400



"\frac{1}{3}<\frac{1}{4}"


"\frac{4}{7}>\frac{2}{5}"


"\frac{2}{8}=\frac{1}{4}"


"\frac{2}{5}<\frac{1}{2}"


"\frac{4}{8}<\frac{3}{4}"


"\frac{4}{5}<1"